Suppose that tanθ=2/y. Find an expression in y for secθ.
Answer by stanbon(75887)

Suppose that tanθ=2/y. Find an expression in y for secθ.
tan(t) = opp/adj
So opp = 2 and adj = y
------------
then r = sqrt(opp^2 + adj^2) = sqrt(2^2+y^2] = sqrt(4+y^2)
-----------------------
sec(t) = r/adj = sqrt(4+y^2)/y
--------------
Cheers,
Stan H.
------------